# Norming Markushevich bases: recent results and open problems

### v1

@@ 5. Banach spaces of continuous functions
Our plan for the section is the following. At first, we shall explain why consideration of the above problem naturally leads to consider the compact space $[0,\omega_{1}]$. Secondly, we describe an important result due to Alexandrov and Plichko [3] that $\mathcal{C}([0,\omega_{1}])$ does not admit a norming M-basis and its recent extension [85]. Finally, we state and comment some problems whose positive answer would answer Problem 5.1 in the positive.
There is an obvious reason why one might think at $[0,\omega_{1}]$ when considering Problem 5.1, namely the fact that $[0,\omega_{1}]$ is a rather simple and well studied example of a scattered compact space that is not Eberlein. However, there is much more than that. Consider the particular case when the M-basis is actually $1$-norming. Then the evaluation map $\mathsf{ev}$ from (1.1) shows that $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia, thus $\mathcal{K}$ is Valdivia as well, because of [61, Theorem 5.3] and the fact that $\mathcal{K}$ has a dense set of $G_{\delta}$ points (as it is scattered).
Now the main point comes, due to the following important results. First, Deville and Godefroy [22] proved the elegant characterisation that a Valdivia compactum $\mathcal{K}$ is Corson if and only if it doesn’t contain $[0,\omega_{1}]$. Next, by a result of Alster [4] scattered Corson compacta are Eberlein. Therefore, one only has to consider these $\mathcal{K}$ that contain $[0,\omega_{1}]$. For this reason, $[0,\omega_{1}]$ could be considered to be a good test case for Problem 5.1.
@@
$f_{\alpha}\coloneqq\text{1}_{[0,\alpha]}\qquad\mbox{and}\qquad\mu_{\alpha}\coloneqq\delta_{\alpha}-\delta_{\alpha+1}.$
@@
On the other hand, a similar argument only shows that the above system in $\mathcal{C}_{0}([0,\omega_{1}])$ is countably $1/2$-norming and Kalenda [60] proved that $\mathcal{C}_{0}([0,\omega_{1}])$ doesn’t admit any countably $1$-norming M-basis.
Remark 5.4.
It is not hard to check explicitly that the M-basis from Example 5.2 is even a long Schauder basis, see, e.g., [43, Proposition 4.8]. Therefore, differently from the separable case, long Schauder bases might fail to be norming. An analogue example in $\mathcal{C}([0,\omega_{2}])$ shows that long Schauder bases might also fail to be countably norming, [62] (see the discussion concerning $\mathcal{C}([0,\omega_{2}])$ at the end of Section 6).
@@
By definition, every long Schauder basis is a strong M-basis, therefore, we have the existence of a Banach space with strong M-basis, but no norming M-basis. On the other hand, every Banach space with norming M-basis admits a strong M-basis, [3, Theorem 1]; this is actually true more generally for every Plichko space and even for spaces with a projectional skeleton [65, Theorem 1.2] and it is a standard transfinite induction argument from Terenzi’s result [93] for separable spaces.
Instead of giving the full proof of Theorem 5.5, we will follow [2] and prove that the analogue in $\mathcal{C}([0,\eta])$ of the M-basis from Example 5.2 is not norming provided that $\eta\geqslant\omega^{2}$. The idea is to consider a certain ‘ziqqurat like’ function, see (5.1) below, which is also present in [43, Theorem 5.25], [62], and [85]. The proof of Theorem 5.5 then consists in several reductions and stabilisation arguments to show that one can reproduce a function as (5.1), starting from any M-basis of $\mathcal{C}([0,\omega_{1}])$.
@@
[…]
@@
Let us focus on the good news first. An important class for which we can answer Problem 5.1 is the class of compact trees. Trees are an interesting class of compacta because, on the one hand they are a well-behaved generalisation of ordinals, on the other one this bigger flexibility already allows for some interesting examples. As a witness, let us mention some recent results in topology [79, 84, 89, 90] and the celebrated Haydon’s results in renorming [47, 48, 49]. Here we adopt the set-theoretic definition of tree as a poset where the set of predecessors of each element is well-ordered. There are several natural topologies that one can consider on a tree, leading to rather different behaviours, see, e.g., [78]. For our purposes, the correct choice is the so-called coarse-wedge topology (while in renorming theory, as in Haydon’s papers, the interval topology is the suitable one).
@@
It is not difficult to see that if a tree $\mathcal{T}$ has height at most $\omega_{1}$, then it is Corson (because every element of the tree actually has countable height), [79, Theorem 2.8]. On the other hand, if the height of $\mathcal{T}$ is at least $\omega_{1}+1$, we can take an element $t$ with height $\omega_{1}$ and the map $s\mapsto s\wedge t$ is a continuous surjection from $\mathcal{T}$ onto $[0,\omega_{1}]$ (we refer to [85, Section 5] for further details). Consequently, Problem 5.1 has a positive answer when $\mathcal{K}$ is a tree with the coarse-wedge topology. Incidentally, a positive answer was also given for adequate compacta in [45, Section 5].
@@
Now, to the bad news. Problem 5.1 would be solved for all Valdivia compacta if all scattered Valdivia compacta that contain $[0,\omega_{1}]$ admitted a continuous surjection onto it. However, [85, Theorem 4.3] claims that a scattered Valdivia compactum $\mathcal{K}$ admits a continuous image onto $[0,\omega_{1}]$ if and only if it admits a P-point. As it turns out, the compactum $\mathcal{K}_{\varrho}$ that we described in Section 4 fails to admit a P-point and therefore doesn’t admit continuous surjections onto $[0,\omega_{1}]$ (we will give the definition of P-point and the justification of this claim in Section 7). This also explains why Problem 4.13 is open and particularly relevant.
@@
Incidentally, the scattered Valdivia compacta that contain $[0,\omega_{1}]$, but don’t admit continuous surjections onto it must necessarily contain weak P-points and fail to contain P-points, see [85, Remark 4.7]. Therefore, they have to resemble the properties of $\mathcal{K}_{\varrho}$, which once more hints at the importance of such a construction for Godefroy’s problem.
@@
As we saw, the methods from [3, 85] alone aren’t efficient enough to yield a positive answer to Problem 5.1. Hence, to conclude our discussion of the problem we shall highlight two problems whose positive answer would solve the main question in the positive, in combination with [85]. The first one concerns the assumption on $\mathcal{K}$ to be Valdivia that we made above.
@@
Recall that by [61, Theorem 5.3], this would be equivalent to the fact that $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia. For a general Banach space it is known that countably norming M-basis does not imply Valdivia dual ball; in fact, in the already mentioned [60] Kalenda actually proved that the dual ball of $\mathcal{C}_{0}([0,\omega_{1}])$ is not Valdivia. However, here we are only asking about $\mathcal{C}(\mathcal{K})$ spaces and the main concern is on norming M-bases.
@@
A positive answer to both Problem 5.9 and Problem 5.10 would answer in the positive Godefroy’s problem for $\mathcal{C}(\mathcal{K})$ spaces, Problem 5.1. Indeed, by the positive answer to the first problem, one can assume that $\mathcal{K}$ is Valdivia; then one has to distinguish two cases. If $\mathcal{K}$ doesn’t contain $[0,\omega_{1}]$, it is Eberlein by Alster’s [4] and Deville–Godefroy’s [22] results that we mentioned previously in the section. If $\mathcal{K}$ contains $[0,\omega_{1}]$, the positive answer to the second problem contradicts Theorem 5.8.
@@
Both examples depend on the fact that if $\mathcal{K}$ is c.c.c. (i.e., if every collection of mutually disjoint non-empty open subsets of $\mathcal{K}$ is countable), then $\mathcal{C}([0,\omega_{1}])$ doesn’t embed in $\mathcal{C}(\mathcal{K})$. Indeed, $\mathcal{K}$ being c.c.c. is equivalent to $\mathcal{C}(\mathcal{K})$ not containing $c_{0}(\omega_{1})$, [31, Theorem 14.26], while $\mathcal{C}([0,\omega_{1}])$ plainly does contain it.
For the first example it is then enough to take the cube $\{0,1\}^{\omega_{1}}$, which is a Valdivia compactum containing $[0,\omega_{1}]$. Moreover, a standard application of the Delta system lemma shows that $\{0,1\}^{\omega_{1}}$ is c.c.c.. Alternatively, $\mathcal{C}([0,\omega_{1}])$ doesn’t embed in $\mathcal{C}(\{0,1\}^{\omega_{1}})$, because the latter has a norming M-basis, by [45, Theorem 5.1].
The second example is less elementary and it requires the existence of a compactification $\mathcal{K}=\gamma\omega$ of $\omega$ whose remainder $\gamma\omega\setminus\omega$ is $[0,\omega_{1}]$. This is an example from [34], based on Magill’s theorem [72, Theorem 2.1]; see [102, pp. 141–145] for details. Then it is clear that $\mathcal{K}$ is scattered and separable, hence c.c.c..

### v2

@@ 4. Norming M-bases and WCG spaces
We now enter the last part of the section, where we show that Godefroy’s problem, Problem 4.4, has a positive answer if we strengthen the Asplund assumption by asking that $\mathcal{X}$ has countable Szlenk index. The validity of this was suggested to us by Gilles Godefroy and to the best of our knowledge the result has not appeared elsewhere.
Recall that for a $w^{*}$-compact subset $\mathcal{K}$ of $\mathcal{X}^{*}$ and $\varepsilon>0$, $s_{\varepsilon}(\mathcal{K})$ is the set obtained by removing from $\mathcal{K}$ all $w^{*}$-open subsets of $\mathcal{K}$ with diameter less than $\varepsilon$. Then the derivation $s_{\varepsilon}^{\alpha}(\mathcal{K})$ is defined by $s_{\varepsilon}^{\alpha+1}(\mathcal{K})=s_{\varepsilon}(s_{\varepsilon}^{\alpha}(\mathcal{K}))$ and $s_{\varepsilon}^{\alpha}(\mathcal{K})=\bigcap_{\beta<\alpha}s_{\varepsilon}^{\beta}(\mathcal{K})$ for $\alpha$ a limit ordinal. If $s_{\varepsilon}^{\alpha}(\mathcal{K})=\emptyset$ for some $\alpha$, one lets ${\rm Sz}(\mathcal{K},\varepsilon)$ be the least such an ordinal $\alpha$; otherwise, ${\rm Sz}(\mathcal{K},\varepsilon)=\infty$. Finally, the Szlenk index of $\mathcal{K}$ is ${\rm Sz}(\mathcal{K})\coloneqq\sup_{\varepsilon>0}{\rm Sz}(\mathcal{K},\varepsilon)$ and the Szlenk index of $\mathcal{X}$ is ${\rm Sz}(\mathcal{X})\coloneqq{\rm Sz}(B_{\mathcal{X}^{*}})$. Clearly, if $\mathcal{L}\subseteq\mathcal{K}$ are $w^{*}$-compact $s_{\varepsilon}^{\alpha}(\mathcal{L})\subseteq s_{\varepsilon}^{\alpha}(\mathcal{K})$; thus ${\rm Sz}(\mathcal{L})\leqslant{\rm Sz}(\mathcal{K})$. Moreover, it is known that ${\rm Sz}(\mathcal{X})<\infty$ if and only if $\mathcal{X}$ is an Asplund space, [28, Theorem I.5.2]. Therefore, admitting countable Szlenk index is stronger than being Asplund (notice that for an Asplund space $\mathcal{X}$ of density $\kappa$ one has ${\rm Sz}(\mathcal{X})<\kappa^{+}$, [50, Lemma 2.41]). For further information on the Szlenk index we refer to [50, Chapter 2], [80], or Causey’s work, e.g., [14, 15, 16, 17, 18].
The main ingredient in the proof is the following generalisation of a result due to Deville and Godefroy [27] that we will also mention several times in the next section. Further generalisations of Deville–Godefroy’s result can be found in Kalenda’s papers [68, 69], [70, Proposition 3.12].
Proposition 4.14.
Let $\mathcal{X}$ be a Banach space and $\mathcal{K}$ be a $w^{*}$-compact subset of $\mathcal{X}^{*}$. If $\mathcal{K}$ is Valdivia and not Corson, then $\mathcal{K}$ contains a homeomorphic copy of $[0,\omega_{1}]$ that is $\varepsilon$-separated (in norm), for some $\varepsilon>0$.
Notice that the assumption that $\mathcal{K}\subseteq(\mathcal{X}^{*},w^{*})$ does not affect the generality, as every compact set $\mathcal{K}$ is homeomorphic to a $w^{*}$-compact subset of $\mathcal{C}(\mathcal{K})^{*}$.
Proof.
By the result of Deville and Godefroy [27] there exists a homeomorphic embedding $\varphi\colon[0,\omega_{1}]\to(\mathcal{K},w^{*})$. We first notice that the (disjoint) $w^{*}$-compact sets $\varphi([0,\alpha])$ and $\varphi([\alpha+1,\omega_{1}])$ have positive distance (in the norm of $\mathcal{X}^{*}$), for all $\alpha<\omega_{1}$. Indeed, if the distance were 00, there would be sequences $(p_{n})_{n=1}^{\infty}$ in $\varphi([0,\alpha])$ and $(q_{n})_{n=1}^{\infty}$ in $\varphi([\alpha+1,\omega_{1}])$ with $\|p_{n}-q_{n}\|<1/n$. Up to a subsequence, we can assume that $p_{n}$ is $w^{*}$-convergent to some $p\in\varphi([0,\alpha])$. Thus $q_{n}$ $w^{*}$-converges to $p$ as well, which gives the contradiction that $p\in\varphi([\alpha+1,\omega_{1}])$.
Therefore, for $\alpha<\omega_{1}$ we can define
$d_{\alpha}\coloneqq{\rm dist}\big{(}\varphi([0,\alpha]),\varphi([\alpha+1,\omega_{1}])\big{)}>0.$
Hence there are an uncountable subset $\Lambda$ of $[0,\omega_{1}]$ and $\varepsilon>0$ such that $d_{\alpha}\geqslant\varepsilon$ for all $\alpha\in\Lambda$. Moreover, $\overline{\Lambda}$ is homeomorphic to $[0,\omega_{1}]$ and $\Lambda$ is obviously dense in $\overline{\Lambda}$. Thus, up to replacing $\varphi$ with $\varphi\mathord{\upharpoonright}_{\overline{\Lambda}}$ we can assume that $\Lambda$ is dense in $[0,\omega_{1}]$. In particular, all successor ordinals belong to $\Lambda$. Therefore, if $\alpha<\beta\in[0,\omega_{1}]$ are such that there is a successor ordinal $\gamma$ with $\alpha\leqslant\gamma<\beta$, then $\varphi(\alpha)\in\varphi([0,\gamma])$ and $\varphi(\beta)\in\varphi([\gamma+1,\omega_{1}])$, whence $\|\varphi(\alpha)-\varphi(\beta)\|\geqslant d_{\gamma}\geqslant\varepsilon$. In other words, if $\Gamma\coloneqq[0,\omega_{1}]\setminus\{\alpha+1\colon\alpha\mbox{ limit ordinal}\}$, then $\Gamma$ is homeomorphic to $[0,\omega_{1}]$ and $\varphi(\Gamma)$ is $\varepsilon$-separated, as desired. ∎
Remark 4.15.
The above proof actually shows that if $[0,\omega_{1}]\subseteq\mathcal{X}^{*}$, then there is a $w^{*}$-compact subset $\mathcal{L}$ of $[0,\omega_{1}]$ that is homeomorphic to $[0,\omega_{1}]$ and $\varepsilon$-separated. However, this also follows from the statement of the proposition by choosing $\mathcal{K}=[0,\omega_{1}]$.
We are now ready for the proof of the claimed result. As it turns out, we can actually weaken the assumption on the norming M-basis to the one that $\mathcal{X}$ is Plichko.
Theorem 4.16.
Every Plichko space with countable Szlenk index is WCG.
Proof.
Being WCG and having countable Szlenk index are isomorphic properties, so there is no loss in generality in assuming that $\mathcal{X}$ is $1$-Plichko. Hence $(B_{\mathcal{X}^{*}},w^{*})$ is Valdivia. If it is additionally Corson, then $\mathcal{X}$ is WLD, thus WCG by Theorem 4.3. Therefore, we assume towards a contradiction that $(B_{\mathcal{X}^{*}},w^{*})$ is not Corson and we apply Proposition 4.14. Hence, there is a $w^{*}$-compact set $\mathcal{K}\subseteq(B_{\mathcal{X}^{*}},w^{*})$, homeomorphic to $[0,\omega_{1}]$ and that is $\varepsilon$-separated, for some $\varepsilon>0$. For an $\varepsilon$-separated set $\mathcal{L}$ it is clear that $s_{\varepsilon}(\mathcal{L})$ coincides with the set of accumulation points $\mathcal{L}^{\prime}$ of $\mathcal{L}$; therefore, the Szlenk $\varepsilon$-derivation of $\mathcal{K}$ is the same as the Cantor–Bendixson one, $s_{\varepsilon}^{\alpha}(\mathcal{K})=\mathcal{K}^{(\alpha)}$ for every ordinal $\alpha$. Thus we have
$s_{\varepsilon}^{\omega_{1}}(B_{\mathcal{X}^{*}})\geqslant s_{\varepsilon}^{\omega_{1}}(\mathcal{K})=\mathcal{K}^{(\omega_{1})}\neq\emptyset,$
which contradicts the assumption that $\mathcal{X}$ has countable Szlenk index. ∎
@@ 5. Banach spaces of continuous functions
Our plan for the section is the following. At first, we shall explain why consideration of the above problem naturally leads one to consider the compact space $[0,\omega_{1}]$. Secondly, we describe an important result due to Alexandrov and Plichko [3] that $\mathcal{C}([0,\omega_{1}])$ does not admit a norming M-basis and its recent extension [96]. Finally, we state and comment some problems whose positive answers would solve Problem 5.1 in the positive.
There is an obvious reason why one might think of $[0,\omega_{1}]$ when considering Problem 5.1, namely the fact that $[0,\omega_{1}]$ is a rather simple and well studied example of a scattered compact space that is not Eberlein. However, there is much more than that. Consider the particular case of Problem 5.1 when the M-basis is actually $1$-norming. Then the evaluation map $\mathsf{ev}$ from (1.1) shows that $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia, thus $\mathcal{K}$ is Valdivia as well, because of [70, Theorem 5.3] and the fact that $\mathcal{K}$ has a dense set of $G_{\delta}$ points (as it is scattered).
Now the main point comes, due to the following important results. First, Deville and Godefroy [27] proved the elegant characterisation that a Valdivia compactum $\mathcal{K}$ is Corson if and only if it does not contain $[0,\omega_{1}]$. Next, by a result of Alster [4] scattered Corson compacta are Eberlein. Therefore scattered Valdivia compacta that do not contain $[0,\omega_{1}]$ are Eberlein and one only has to consider these $\mathcal{K}$ that contain $[0,\omega_{1}]$. For this reason, $[0,\omega_{1}]$ could be considered to be a good test case for Problem 5.1.
@@
$f_{\alpha}\coloneqq\text{1}_{[0,\alpha]}\qquad\mbox{and}\qquad\mu_{\alpha}\coloneqq\delta_{\alpha}-\delta_{\alpha+1}\qquad(\alpha<\omega_{1}).$
@@
On the other hand, a similar argument only shows that the above system in $\mathcal{C}_{0}([0,\omega_{1}])$ is countably $1/2$-norming and Kalenda [67] proved that $\mathcal{C}_{0}([0,\omega_{1}])$ does not admit any countably $1$-norming M-basis.
@@
It is not hard to check explicitly that the M-basis from Example 5.2 is even a long Schauder basis, see, e.g., [50, Proposition 4.8]. Therefore, unlike the separable case, long Schauder bases might fail to be norming. An analogue example in $\mathcal{C}([0,\omega_{2}])$ shows that long Schauder bases might also fail to be countably norming, [71] (see the discussion concerning $\mathcal{C}([0,\omega_{2}])$ at the end of Section 6).
Remark 5.6.
By definition, every long Schauder basis is a strong M-basis, therefore we have the existence of a Banach space with strong M-basis, but no norming M-basis. On the other hand, every Banach space with norming M-basis admits a strong M-basis, [3, Theorem 1]; this is actually true more generally for every Plichko space and even for spaces with a projectional skeleton [74, Theorem 1.2] and it is a standard transfinite induction argument from Terenzi’s result [104] for separable spaces.
Instead of giving the full proof of Theorem 5.4, we will follow [2] and prove that the analogue in $\mathcal{C}([0,\eta])$ of the M-basis from Example 5.2 is not norming provided that $\eta\geqslant\omega^{2}$. The idea is to consider a certain ‘ziqqurat like’ function, see (5.1) below, which is also present in [50, Theorem 5.25], [71], and [96]. The proof of Theorem 5.4 then consists of several reductions and stabilisation arguments to show that one can reproduce a function as (5.1) starting from any M-basis of $\mathcal{C}([0,\omega_{1}])$.
@@
$f_{\alpha}\coloneqq\text{1}_{[0,\alpha]}\qquad\mbox{and}\qquad\mu_{\alpha}\coloneqq\delta_{\alpha}-\delta_{\alpha+1}$
@@
Let us focus on the good news first. An important class for which we can answer Problem 5.1 is the class of compact trees. Trees are an interesting class of compacta because on the one hand they are a well-behaved generalisation of ordinals and on the other hand this bigger flexibility already allows for several interesting examples. As instances of this let us mention some recent results in topology [90, 95, 100, 101] and Haydon’s celebrated results in renorming [54, 55, 56]. Here we adopt the set-theoretic definition of tree as a poset where the set of predecessors of each element is well-ordered. There are several natural topologies that one can consider on a tree, leading to rather different behaviours, see, e.g., [89]. For our purposes, the correct choice is the so-called coarse-wedge topology (while in renorming theory, as in Haydon’s papers, the interval topology is the suitable one).
@@
It is not difficult to see that if a tree $\mathcal{T}$ has height at most $\omega_{1}$, then it is Corson (because every element of the tree actually has countable height), [90, Theorem 2.8]. On the other hand, if the height of $\mathcal{T}$ is at least $\omega_{1}+1$, we can take an element $t$ with height $\omega_{1}$ and the map $s\mapsto s\wedge t$ is a continuous surjection from $\mathcal{T}$ onto $[0,\omega_{1}]$ (we refer to [96, Section 5] for further details). Consequently, Problem 5.1 has a positive answer when $\mathcal{K}$ is a tree with the coarse-wedge topology. Incidentally, a positive answer was also given for adequate compacta in [52, Section 5].
@@
Now, to the bad news. Problem 5.1 would be solved for all Valdivia compacta if all scattered Valdivia compacta that contain $[0,\omega_{1}]$ admitted a continuous surjection onto it. However, [96, Theorem 4.3] claims that, for a scattered Valdivia compactum $\mathcal{K}$, there is a continuous map from $\mathcal{K}$ onto $[0,\omega_{1}]$ if and only if $\mathcal{K}$ admits a P-point. As it turns out, the compactum $\mathcal{K}_{\varrho}$ that we described in Section 4 fails to admit a P-point and therefore does not admit continuous surjections onto $[0,\omega_{1}]$ (we will give the definition of P-point and the justification of this claim in Section 7). This also explains why Problem 4.13 is open and particularly relevant.
@@
Incidentally, the scattered Valdivia compacta that contain $[0,\omega_{1}]$, but do not admit continuous surjections onto $[0,\omega_{1}]$ must necessarily contain weak P-points and fail to contain P-points, see [96, Remark 4.7]. Therefore, they have to resemble the space $\mathcal{K}_{\varrho}$, which once more hints at the importance of such a construction for Godefroy’s problem.
@@
As we saw, the methods from [3, 96] alone are not efficient enough to yield a positive answer to Problem 5.1 in its full generality. Hence, to conclude our discussion of the problem we shall highlight two problems whose positive answer would solve the main question in the positive, in combination with [96]. The first one concerns the assumption on $\mathcal{K}$ to be Valdivia that we made above.
@@
Recall that by [70, Theorem 5.3], this would be equivalent to the fact that $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia. For a general Banach space it is known that the existence of a countably norming M-basis does not imply that the dual ball is Valdivia; in fact, in the already mentioned [67] Kalenda actually proved that the dual ball of $\mathcal{C}_{0}([0,\omega_{1}])$ is not Valdivia. However, here we are only asking about $\mathcal{C}(\mathcal{K})$ spaces and the main concern is on norming M-bases.
@@
A positive answer to both Problem 5.9 and Problem 5.10 would answer in the positive Godefroy’s problem for $\mathcal{C}(\mathcal{K})$ spaces, Problem 5.1. Indeed, by the positive answer to the first problem, one can assume that $\mathcal{K}$ is Valdivia; then one has to distinguish two cases. If $\mathcal{K}$ does not contain $[0,\omega_{1}]$, it is Eberlein by Alster’s [4] and Deville–Godefroy’s [27] results that we mentioned previously in the section. If $\mathcal{K}$ contains $[0,\omega_{1}]$, the positive answer to the second problem contradicts Theorem 5.8.
@@
Both examples depend on the fact that if $\mathcal{K}$ is c.c.c. (i.e., if every collection of mutually disjoint non-empty open subsets of $\mathcal{K}$ is countable), then $\mathcal{C}([0,\omega_{1}])$ does not embed in $\mathcal{C}(\mathcal{K})$. Indeed, $\mathcal{K}$ being c.c.c. is equivalent to $\mathcal{C}(\mathcal{K})$ not containing $c_{0}(\omega_{1})$, [38, Theorem 14.26], while $\mathcal{C}([0,\omega_{1}])$ plainly does contain it (for instance, as $[0,\omega_{1}]$ is not c.c.c.).
For the first example it is then enough to take the cube $\{0,1\}^{\omega_{1}}$, which is a Valdivia compactum containing $[0,\omega_{1}]$. Moreover, a standard application of the Delta system lemma shows that $\{0,1\}^{\omega_{1}}$ is c.c.c.. Alternatively, $\mathcal{C}([0,\omega_{1}])$ does not embed in $\mathcal{C}(\{0,1\}^{\omega_{1}})$, because the latter has a norming M-basis, by [52, Theorem 5.1].
The second example is less elementary and it requires the existence of a compactification $\mathcal{K}=\gamma\omega$ of $\omega$ whose remainder $\gamma\omega\setminus\omega$ is $[0,\omega_{1}]$. This is an example from [41], based on Magill’s theorem [82, Theorem 2.1]; see [113, pp. 141–145] for details. Then it is clear that $\mathcal{K}$ is scattered and separable, hence c.c.c..